Grade 12Physical Chemistry

Q. 25.3 g of sodium carbonate is dissolved in enough water to make 250 mL of solution.If sodium carbonate dissociates completely,molar concentration of sodium ion,Na+ and carbonate ions,CO3^2- are respectively..????

Aditi
5 Years agoGrade 12

1 Answer

Pawan Prajapati
5 Years ago
Answer :

Moles of Na2CO3 = mass/molar mass
= 25.3/106
= 0.239 moles
Molarity of Na2CO3 solution = number of moles/ volume in lire
= 0.239/0.25 (250ml = 0.25 L)
= 0.955 M
Now, 1 mole of Na2CO3 contains 2 moles of Na+ and 1 mole of CO32- ions.
Therefore,
Concentration of Na+ = 2 x 0.955M = 1.91M
Concentration of CO32- = 1 x 0.955M = 0.955M

Thanks